Hi! I'm not getting one thing on my math homework, can you help me please?
OK...here's the word problem.

- Dough bought a small tree at the local garden center to plant in his backyard. The employee at the garden center told him that the tree would grow two centimeters per week. The linear equation below represents, h the height of the tree in Doug's backyard. H=2c+61
-What I'm not getting is, what does 61 represent. How tall the tree is when he got it?

Found 2 solutions by jim_thompson5910, fcabanski:
Answer by jim_thompson5910(35256)

When c = 0, ie when he first got it, H = 2(0)+61 = 61 which means that the height when he got it was 61 cm


So the 61 represents the initial or starting height of the plant.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------

Answer by fcabanski(1391)

61 represents the initial height, when the number of weeks after planting is 0.


Every linear equation has a y-intercept (in this case h-intercept). That is the value of the equation when the independent variable, usually x but in this case c, is 0.


When c=0 then this equation is h=2*0+61 = 61.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com